Find the nullity and the range of the space spanned by ๐‘ข = (1,0,2), ๐‘ฃ = (โˆ’3,1,1), ๐‘ค = (0,1,4) and ๐‘ฅ = (โˆ’1, โˆ’3,5)


1
Expert's answer
2022-03-05T07:38:08-0500

"\begin{bmatrix}
 1&-3&0&-1\\
 0&1&1&-3\\2&1&4&5
\end{bmatrix}"





"R_3-2R_1\to\>R_3"

"R_1+3R_2\to\>R_1"


"\begin{pmatrix}
 1&&0&&3& &-10\\
 0&&1&&1&&-3\\
0&&7&&4&&7
\end{pmatrix}"



"\frac{R_3-7R_2}{-3}\to\>R_3"



"\begin{pmatrix}
 1&&0&& 3&&-10 \\
 0&&1& & 1&&-3\\
0&&0&&1&&\frac{-28}{3}
\end{pmatrix}"




"R_1-3R_3\to\>R_1"

"R_2-R_3\to\>R_2"



rref is "\begin{bmatrix}
 1&0&0& 18 \\
 0&1&0&\frac{19}{3}\\0&0&1&-\frac{28}{3}
\end{bmatrix}"



Nullity =1


Range="[\begin{bmatrix}
 1 \\0\\
 2
\end{bmatrix},\begin{bmatrix}
 -3 \\1\\
 1
\end{bmatrix},\begin{bmatrix}
 0\\
 1\\4
\end{bmatrix}]"